# A Measure Based Generalizable Approach to Understandability

### v1

@@ 1. Introduction
As a starting point, this paper focuses on one kind of human priors and sensibilities, namely understandability. We propose a cognitive-science-rooted framework to organize fundamental aspects of understandability into six coherent dimensions: perceptual advantage, memory cost, pattern decodability, cohesion, logical consistency and semantic fit. While existing research on understandability measures is fragmented, we survey various such efforts across domains from the perspective of this framework. This should highlight research gaps and motivate future conversations for developing more generalized measures that can work “as a whole” across a variety of information entities (text, code, etc.).
@@ 2.1. Cognitive modeling
The field of cognitive science concerns with understanding processes in the human mind, examples are attention, speech, comprehension and learning. One way in which psychologists study these phenomena is by: 1) developing specific hypotheses about the process under study, 2) developing computational models and architectures that implement these hypotheses and 3) evaluating the predictions of the models against actual human behavior for a specific task. Some examples are (kintsch2013construction, ; xu2005cognitive, ; oudeyer2007intrinsic, ; mervis1981categorization, ). A longstanding criticism of task specific cognitive models has been that they lead to a fragmented approach and make cumulative progress difficult (laird2017standard, ). As a solution, calls have been made for integrated models of cognition. Cognitive architectures (e.g., ACT-R, Soar, Sigma) (laird2017standard, ) are a step in fulfilling that goal. “However, there has historically been little agreement either across or within specialties as to the overall nature and shape of this architecture” (laird2017standard, ). The larger goal of unified theories of cognition (newell1994unified, ) remains a work in progress (laird2017standard, ).
@@ 2.2. Cognitive models of understanding
Understanding is a specific kind of cognitive process and similar issues befall it. While there is a general idea that 1) understanding involves some sort of fit between data or the artifact being understood with the frame or schema (e.g., a mental structure for organizing and perceiving new information) (newell1994unified, ; russell1993cost, ; naumer2008sense, ) and 2) that it involves activation and evaluation of relevant information in memory (marinier2006cognitive, ), more detailed models of comprehension are fragmented and domain-specific. For example, there are models of text comprehension (mcnamara2009toward, ), program comprehension (xu2005cognitive, ), language comprehension (johnson1981comprehension, ) and so on. With this in mind, we make an attempt to describe major aspects of the human process of understanding, based on emerging ideas. Our main aim here is to figure out the key aspects of the underlying process of understanding and delineating what aspects of an artifact (e.g., text, code) contribute to their understandability. This gives us an initial base of understandability aspects to be investigated and improved upon in future.
@@
Whereas mechanisms involved in perception, memory (retrieval and updation) are largely task independent, evaluations can be more task dependent (e.g., novelty of stimulus for learning and motivation, matching with goal state for simple planning). For comprehension, models across domains (text comprehension (kintsch2013construction, ), language comprehension (johnson1981comprehension, ), code comprehension (xu2005cognitive, ) and general sense-making models (russell1993cost, ; naumer2008sense, )) largely suggest two kinds of evaluations, namely cohesion and coherence (also referred to as data-schema fit (mekler2019framework, )). Cohesion is about making sense of a structure through its individual constituents (e.g., words in a sentence, methods in a program class), while coherence is about relation between multiple structures (e.g., how paragraphs make an argument, logical arrangement of UI screens, relations of a program class with other classes). In terms of measures, cohesion is often characterized as relatedness or similarity between the constituents of a structure (e.g., use of the same word, pronouns) while coherence has two parts when it comes to comprehension, namely logical consistency and quality of semantic relationships (thagard2002coherence, ) (or “semantic fit”). Finally, the number of cognitive cycles is also a measure of the complexity of the artifact being understood; we can operationalize it as some measure of difficulty of decoding a given pattern in an artifact (or “pattern decodability”).
@@ 3. Organizing understandability into dimensions
Summarizing the emerging landscape across cognitive models of understanding, we propose six dimensions of understandability: perceptual advantage, memory cost, pattern decodability, cohesion, logical consistency and semantic fit. We highlight these dimensions with underlines in Figure  1 that summarizes the cognitive processes of understanding, synthesized across various recent models (namely, cognitive architectures, models of comprehension and sensemaking). The rooting of these models in cognitive architectures allows for coherence, while the use of abstractions from cognitive psychology allows for generalizability across domains.
@@ 3.3. Pattern decodability
Notion of computational depth (antunes2006computational, ) adds on top of above measures characterizing the existence of patterns. It measures the amount of non-trivial patterns in a string as a difference between time restricted Kolmogorov complexity and its traditional counterpart (antunes2006computational, ). A string having a short compressed version but requiring considerable computation time to uncompress will have large computational depth, compared to random strings or strings with trivial patterns. Further, there are computable variants of computation depth such as Lempel-Ziv depth (jordon2023pushdown, ). If we can represent any information entity as a binary string, the above collection of measures can serve us in quantifying the amount of patterns in the entity and how computationally intensive those patterns are.
@@ 3.5. Logical consistency
Information entities in real life often contain inconsistencies. Humans are able to reason under uncertainty and inconsistent data to a certain extent, but in general, the greater the inconsistency, the lesser the entity makes sense to the reader, and requires more processing (and thus, greater time and cognitive effort) before reasoning can be performed on it. Measures of inconsistency can also be used to evaluate how far a given entity is from satisfying a given set of constraints or goals. These constraints/goals can come from the problem which the human is trying to solve, or from the schema used for making sense of the data.
@@ 4. Concluding remarks and future directions
The notion of understandability has been operationalized in various domains in fragmented ways, with several measures and rules of thumb existing to make content more understandable to their readers (e.g., usability heuristics for UI design, metrics for clean code, text comprehensibility measures used in psychology experiments). This fragmentation not only hurts building a coherent body of knowledge in HCI, but is also a practical limitation, such as when building agents that can be steered by humans in generating understandable content for various tasks (e.g., general-purpose LLMs).
However, such efforts in various domains are meant to deal with fundamental human cognitive abilities and limitations, and thus in this paper, we aimed to derive six dimensions of understandability from widely accepted cognitive models of how humans go about understanding “stuff”. Our six dimensions, namely perceptual advantage, memory cost, pattern decodability, cohesion, logical consistency, and semantic fit, each map on to various stages of the cognitive processes underlying understanding, and the existing measures of understandability map on to these dimensions.
@@
However, despite solid theoretical grounding, we point out that our effort is preliminary, and, as with any theory-building effort, our six dimensions of understandability need validation. Thus, an essential, immediate future work is to critically evaluate the completeness, appropriateness and generalizability of our dimensions via user studies. In turn, this will also serve as indirect validation of the cognitive models in terms of their generalizability and completeness.
@@
Finally, going beyond understandability and information processing, this paper offers preliminary insights into a method for drawing from theories of cognitive psychology to make them practical to HCI research. For example, there are multiple kinds of human priors and sensibilities that are relevant to HCI (e.g., user motivation (schmidhuber2010formal, ; oudeyer2007intrinsic, ), exploration tendencies (pirolli1995information, )). They deserve research conversations of their own, but can take a similar approach to us in organizing and exploring those aspects with a solid grounding in theories.

### v2

@@ 1. Introduction
As a starting point, this paper focuses on one kind of human priors and sensibilities, namely understandability. We propose a cognitive-science-rooted framework to organize fundamental aspects of understandability into six coherent dimensions: perceptual quality, memory cost, pattern decodability, cohesion, logical consistency and semantic fit. While existing research on understandability measures is fragmented, we survey various such efforts across domains from the perspective of this framework. This should highlight research gaps and motivate future conversations for developing more generalized measures that can work “as a whole” across a variety of information entities (text, code, etc.).
@@ 2.1. Cognitive modeling
The field of cognitive science concerns with understanding processes in the human mind, examples are attention, speech, comprehension and learning. One way in which psychologists study these phenomena is by: 1) developing specific hypotheses about the process under study, 2) developing computational models and architectures that implement these hypotheses and 3) evaluating the predictions of the models against actual human behavior for a specific task. Some examples are (kintsch2013construction, ; von1995program, ; oudeyer2007intrinsic, ; mervis1981categorization, ). A longstanding criticism of task specific cognitive models has been that they lead to a fragmented approach and make cumulative progress difficult (laird2017standard, ). As a solution, calls have been made for integrated models of cognition. Cognitive architectures (e.g., ACT-R, Soar, Sigma) (laird2017standard, ) are a step in fulfilling that goal. “However, there has historically been little agreement either across or within specialties as to the overall nature and shape of this architecture” (laird2017standard, ). The larger goal of unified theories of cognition (newell1994unified, ) remains a work in progress (laird2017standard, ).
@@ 2.2. Cognitive models of understanding
Understanding is a specific kind of cognitive process and similar issues befall it. While there is a general idea that 1) understanding involves some sort of fit between data or the artifact being understood with the frame or schema (e.g., a mental structure for organizing and perceiving new information) (newell1994unified, ; russell1993cost, ; naumer2008sense, ) and 2) that it involves activation and evaluation of relevant information in memory (marinier2006cognitive, ), more detailed models of comprehension are fragmented and domain-specific. For example, there are models of text comprehension (mcnamara2009toward, ), program comprehension (von1995program, ), language comprehension (johnson1981comprehension, ) and so on. However, while the details of comprehension models may vary both within and across domains, the key aspects are largely shared. With this in mind, we make an attempt to describe major aspects of the human process of understanding, based on emerging ideas. Our main aim here is to figure out the key aspects of the underlying process of understanding and delineating what aspects of an artifact (e.g., text, code) contribute to their understandability. This gives us an initial base of understandability aspects to be investigated and improved upon in future.
@@
Whereas mechanisms involved in perception, memory (retrieval and updation) are largely task independent, evaluations can be more task dependent (e.g., novelty of stimulus for learning and motivation, matching with goal state for simple planning). For comprehension, models across domains (text comprehension (kintsch2013construction, ), language comprehension (johnson1981comprehension, ), code comprehension (von1995program, ) and general sense-making models (russell1993cost, ; naumer2008sense, )) largely suggest two kinds of evaluations, namely cohesion and coherence (also referred to as data-schema fit (mekler2019framework, )). Cohesion is about making sense of a structure through its individual constituents (e.g., words in a sentence, methods in a program class), while coherence is about relation between multiple structures (e.g., how paragraphs make an argument, logical arrangement of UI screens, relations of a program class with other classes). In terms of measures, cohesion is often characterized as relatedness or similarity between the constituents of a structure (e.g., use of the same word, pronouns) while coherence can be seen as a constraint satisfaction process involving varieties of constraints based on the kinds of elements and relationships involved (thagard2002coherence, ). Coherence constraints can be of broadly two main types: 1) logical consistency between declaration or proposition like elements and 2) semantic compatibility or fit between elements in relationships such as analogies, explanations, interpretations and conceptual associations (thagard2002coherence, ). Finally, the number of cognitive cycles is also related to the inherent complexity of the artifact being understood (madl2011timing, ); we can operationalize it as some measure of difficulty of mining useful patterns present in an artifact (or “pattern decodability”).
@@ 3. Organizing understandability into dimensions
Summarizing the emerging landscape across cognitive models of understanding, we propose six dimensions of understandability: perceptual quality, memory cost, pattern decodability, cohesion, logical consistency and semantic fit. We highlight these dimensions with underlines in Figure  1 that summarizes the cognitive processes of understanding, synthesized across various recent models (namely, cognitive architectures, models of comprehension and sensemaking). The rooting of these models in cognitive architectures allows for coherence, while the use of abstractions from cognitive psychology allows for generalizability across domains.
@@ 3.3. Pattern decodability
Notion of computational depth (antunes2006computational, ) adds on top of above measures characterizing the existence of patterns. It measures the amount of non-trivial patterns in a string as a difference between time restricted Kolmogorov complexity and its traditional counterpart (antunes2006computational, ). A string having a short compressed version but requiring considerable computation time to uncompress will have large computational depth, compared to random strings or strings with trivial patterns. Further, there are computable variants of computation depth such as Lempel-Ziv depth (jordon2023pushdown, ). If we can represent any information entity as a binary string, the above collection of measures can serve us in quantifying the amount of useful patterns in the entity and how computationally intensive mining those patterns is.
@@ 3.5. Logical consistency
Information entities in real life often contain inconsistencies between their declaration or proposition like elements. Humans are able to reason under uncertainty and inconsistent data to a certain extent, but in general, the greater the inconsistency, the lesser the entity makes sense to the reader, and requires more processing (and thus, greater time and cognitive effort) before reasoning can be performed on it. Measures of inconsistency can also be used to evaluate how far a given entity is from satisfying a given set of constraints or goals. These constraints/goals can come from the problem which the human is trying to solve, or from the previously learned schema used for making sense of the data.
@@ 4. Concluding remarks and future directions
The notion of understandability has been operationalized in various domains in fragmented ways, with several measures and rules of thumb existing to make content more understandable to their readers (e.g., usability heuristics for UI design, metrics for clean code, text comprehensibility measures used in psychology experiments). This fragmentation not only hurts building a coherent body of knowledge in HCI, but is also a practical limitation, such as when building agents that can be steered by humans in generating understandable content for various tasks.
However, such efforts in various domains are meant to deal with fundamental human cognitive abilities and limitations, and thus in this paper, we aimed to derive six dimensions of understandability from widely accepted cognitive models of how humans go about understanding “stuff”. Our six dimensions, namely perceptual quality, memory cost, pattern decodability, cohesion, logical consistency, and semantic fit, each map on to various stages of the cognitive processes underlying understanding, and the existing measures of understandability map on to these dimensions.
@@
However, despite solid theoretical grounding, we point out that our effort is preliminary, and, as with any theory-building effort, our six dimensions need validation. Thus, an essential, immediate future work is to critically evaluate the completeness, appropriateness and generalizability of our dimensions via user studies. In turn, this will also serve as indirect validation of the cognitive models in terms of their generalizability and completeness.
@@
Finally, going beyond understandability, this paper offers preliminary insights into a method for drawing from theories of cognitive psychology to make them practical to HCI research. For example, there are multiple kinds of human priors and sensibilities that are relevant to HCI (e.g., user motivation (schmidhuber2010formal, ; oudeyer2007intrinsic, ), exploration tendencies (pirolli1995information, )). They deserve research conversations of their own, but can take a similar approach to us in organizing and exploring those aspects with a solid grounding in theories.